# S2AFormer: Strip Self-Attention for Efficient Vision Transformer

### v1

@@ IV-D Qualitative Results
The qualitative results in Fig. 4 highlight the efficacy of our method in generating visually coherent and fine-grained segmentation maps on the ADE20K validation set [79]. Complementing these findings, Figs. 5 and 6 showcase task-specific visualizations from our S2AFormer-S model, addressing instance segmentation and object detection challenges, respectively. These outcomes underscore the model’s robustness in accurately localizing and delineating objects across diverse and cluttered scenes, even under varying scales and occlusions.
@@ IV-E Inference Speed
We implemented S2AFormer across three distinct platforms—GPU, CPU, and ONNX—and evaluated its efficiency through throughput analysis. Table V compares throughput across these platforms, tested on an NVIDIA H100 NVL GPU and an Intel(R) Xeon(R) Gold 6426Y CPU.
As shown in the table, for GPU-based inference, S2AFormer demonstrates near-optimal performance with comparable ImageNet-1K accuracy. For example, S2AFormer-mini achieves a throughput of 6330 images/s on GPU, significantly outperforming other models. It exceeds the throughput of EMO-2M [13], which achieves 3359 images/s, making S2AFormer-mini nearly twice as fast in GPU throughput. This highlights its strong image processing capabilities compared to similar models like EfficientViT-M3 [11] (4355 images/s) and PVT-T [25] (5033 images/s).
In CPU and ONNX-based inference, S2AFormer also shows highly competitive performance. For example, S2AFormer-T surpasses EdgeViT-XS [12] in throughput on both CPU (41.75 images/s) and ONNX (112.89 images/s) platforms. This highlights the model’s robustness in non-GPU environments, where inference speed can become a bottleneck. This is particularly important for real-world deployment scenarios, where model efficiency across various hardware platforms, including CPUs and optimized ONNX models, plays a crucial role in practical applications.
TABLE VII: Performance comparisons with and without LIM on COCO val2017 dataset [81].
Model LIM Top-1$\uparrow$ RetinaNet 1$\times$
@@
TABLE VIII: Performance comparisons between convolution and pooling operations for spatial reduction purposes using Top-1 accuracy and inference throughput.
@@ IV-F1 Effectiveness of LIM
As shown in Table VII, for S2AFormer-Mini, the parameter count increases slightly from 11.55M to 11.65M, while the GFLOPs remain nearly unchanged, with values of 159.06 and 159.20, respectively. Initially, the model achieves a Top-1 accuracy of 74.89%. However, after incorporating LIM, there is a noticeable improvement in performance, with the Top-1 accuracy rising to 75.06%. A similar trend is observed for S2AFormer-T. Without LIM, the model achieves a Top-1 accuracy of 77.63%, but with LIM, the accuracy increases slightly to 77.73%. The parameter count for this model increases modestly from 12.29M to 12.44M, and GFLOPs show a minor increase from 163.50 to 163.71. Despite these small increases in computational cost, the model’s Average Precision (AP) improves significantly, with LIM boosting the AP from 35.3 to 36.7. This demonstrates that the improvements in accuracy come with minimal additional computational overhead, particularly in terms of detection performance.
@@ IV-F2 Effectiveness of Conv Spatial Reduction
As shown in Table VIII, our comparison between Convolution and Pooling for reducing spatial dimensions reveals that Convolution slightly outperforms Pooling in both the S2AFormer-mini and S2AFormer-T models. Specifically, S2AFormer-mini achieves a Top-1 accuracy of 75.1% with Convolution, which is slightly higher than the 75.0% obtained with Pooling. Similarly, S2AFormer-T reaches a Top-1 accuracy of 77.7% using Convolution, compared to 77.3% with Pooling.

### v2

@@ IV-D Qualitative Results
The qualitative results in Fig. 6 highlight the efficacy of our method in generating visually coherent and fine-grained segmentation maps on the ADE20K validation set [zhou2017scene]. Complementing these findings, Figs. 7 and 8 showcase task-specific visualizations from our S2AFormer-S model, addressing instance segmentation and object detection challenges, respectively. These outcomes underscore the model’s robustness in accurately localizing and delineating objects across diverse and cluttered scenes, even under varying scales and occlusions.
TABLE VII: Throughput comparison across GPU, CPU, and ONNX platforms. GPU: NVIDIA H100 NVL, CPU: Intel(R) Xeon(R) Gold 6426Y.
[…]
EMO-1M [zhang2023rethinking] 1.3 71.5 4674 45.57 118.72
EfficientViT-M3 [liu2023efficientvit] 6.9 73.4 4355 35.34 169.35
MobileMamba-T2 [he2024mobilemamba] 8.8 73.6 6279 63.75 178.38
EfficientViT-M4 [liu2023efficientvit] 8.8 74.3 4289 34.34 159.51
EMOv2-1M [zhang2025emov2] 1.4 72.3 4886 45.70 112.98
EMO-2M [zhang2023rethinking] 2.3 75.1 3359 41.20 83.79
PVT-T [wang2021pyramid] 13.2 75.1 5033 61.01 117.63
S2AFormer-mini 5.0 75.1 6330 58.21 118.26
FastViT-T8 [vasu2023fastvit] 3.6 75.6 7615 50.76 129.27
EfficientFormerV2-S0 [li2023rethinking] 3.5 75.7 520 44.57 136.32
MobileMamba-T4 [he2024mobilemamba] 14.2 76.1 5124 52.33 146.37
EdgeViT-XXS [chen2022edgevit] 4.1 74.4 3614 50.82 134.79
PoolFormer-S12 [yu2024metaformer] 12.0 77.2 2148 32.67 102.16
EfficientViT-M5 [liu2023efficientvit] 12.4 77.1 3843 28.48 126.28
EdgeViT-XS [chen2022edgevit] 6.7 77.5 2872 35.14 106.10
S2AFormer-T 5.8 77.7 4768 41.75 112.89
EMO-5M [zhang2023rethinking] 5.1 78.4 2370 38.18 62.52
PVTv2-B1 [wang2021pvtv2] 14.0 78.7 3930 79.51 37.96
EMO-6M [zhang2023rethinking] 6.1 79.0 2102 35.90 57.00
StarNet-S4 [ma2024rewrite] 7.5 78.4 2570 30.83 64.26
EfficientFormerV2S1 [li2023rethinking] 6.1 79.0 472 37.89 104.35
EfficientFormer-L1 [li2022efficientformer] 12.3 79.2 3572 66.93 88.72
iFormer-S [zheng2025iformer] 6.5 78.8 3764 32.89 66.30
S2AFormer-XS 6.5 78.9 4185 34.09 71.63
FastViT-T12 [vasu2023fastvit] 6.8 79.1 5277 42.06 88.31
FastViT-S12 [vasu2023fastvit] 8.8 79.8 4952 39.41 68.09
FastViT-SA12 [vasu2023fastvit] 10.9 80.6 4812 38.48 50.12
LSNet-B [wang2025lsnet] 23.2 80.3 2890 27.20 52.88
PVT-S [wang2021pyramid] 24.5 79.8 2736 43.50 53.01
EdgeViT-S [chen2022edgevit] 11.1 81.0 1951 23.15 58.32
EfficientViT-M5 [liu2023efficientvit] 12.4 80.8 2396 29.05 54.62
S2AFormer-S 10.7 80.8 2644 25.78 55.84
PVT-M [wang2021pyramid] 44.2 81.2 $17$ 25.96 36.89
EfficientFormerV2-S2 [li2023rethinking] 12.6 81.6 256 25.08 46.94
PVT-L [wang2021pyramid] 61.4 81.7 587 18.84 27.77
EfficientVMamba-B [pei2024efficientvmamba] 33.0 81.8 2432 26.75 32.54
PVTv2-B2 [wang2021pvtv2] 25.4 82.0 2218 40.20 40.83
EfficientFormer-L3 [li2022efficientformer] 31.3 82.4 1550 35.96 25.31
VRWKV-B [duan2024vrwkv] 93.7 82.0 1520 20.49 24.68
S2AFormer-M 24.9 82.3 1652 15.50 27.15
@@ IV-E Inference Speed
We implemented S2AFormer on three different platforms and evaluated its efficiency through throughput analysis. Note that, to ensure fairness and reproducibility, we report GPU, CPU, and ONNX throughput using a single codebase and the same checkpoint, rather than referencing results from prior works. All experiments were conducted on an NVIDIA H100 NVL GPU and an Intel Xeon Gold 6426Y CPU, using PyTorch 2.5.1 (for GPU/CPU) and ONNX Runtime (for ONNX).
As shown in Table VII, for GPU-based inference, S2AFormer demonstrates near-optimal performance with comparable ImageNet-1K accuracy. For example, S2AFormer-mini achieves a throughput of 6330 images/s on the GPU, significantly outperforming other models. It exceeds the throughput of EMO-2M [zhang2023rethinking], which achieves 3359 images/s, making S2AFormer-mini nearly twice as fast in GPU throughput. This highlights its strong image processing capabilities compared to similar models like EfficientViT-M3 [liu2023efficientvit] (4355 images/s) and PVT-T [wang2021pyramid] (5033 images/s).
In CPU and ONNX-based inference, S2AFormer also shows highly competitive performance. For example, S2AFormer-T surpasses EdgeViT-XS [chen2022edgevit] in throughput on both CPU (41.75 images/s) and ONNX (112.89 images/s) platforms. This highlights the model’s robustness in non-GPU environments, where inference speed can become a bottleneck. This is particularly important for real-world deployment scenarios, where model efficiency across various hardware platforms, including CPUs and optimized ONNX models, plays a crucial role in practical applications.
IV-F Ablation Studies
TABLE VIII: Performance comparisons with and without LIM on COCO val2017 dataset [lin2014microsoft].
[…]
@@ IV-F Ablation Studies
TABLE IX: Performance comparisons between convolution and pooling operations for spatial reduction purposes using Top-1 accuracy and inference throughput.
@@ IV-F1 Effectiveness of LIM
As shown in Table VIII, for S2AFormer-Mini, the parameter count increases slightly from 11.55M to 11.65M, while the GFLOPs remain nearly unchanged, with values of 159.06 and 159.20, respectively. Initially, the model achieves a Top-1 accuracy of 74.89%. However, after incorporating LIM, there is a noticeable improvement in performance, with the Top-1 accuracy rising to 75.06%. A similar trend is observed for S2AFormer-T. Without LIM, the model achieves a Top-1 accuracy of 77.63%, but with LIM, the accuracy increases slightly to 77.73%. The parameter count for this model increases modestly from 12.29M to 12.44M, and GFLOPs show a minor increase from 163.50 to 163.71. Despite these small increases in computational cost, the model’s Average Precision (AP) improves significantly, with LIM boosting the AP from 35.3 to 36.7. This demonstrates that the improvements in accuracy come with minimal additional computational overhead, particularly in terms of detection performance.
@@ IV-F2 Effectiveness of Conv Spatial Reduction
As shown in Table IX, our comparison between Convolution and Pooling for reducing spatial dimensions reveals that Convolution slightly outperforms Pooling in both the S2AFormer-mini and S2AFormer-T models. Specifically, S2AFormer-mini achieves a Top-1 accuracy of 75.1% with Convolution, which is slightly higher than the 75.0% obtained with Pooling. Similarly, S2AFormer-T reaches a Top-1 accuracy of 77.7% using Convolution, compared to 77.3% with Pooling.
@@
IV-G Limitations and Future Works
While our proposed S2AFormer strikes an effective balance between efficiency and accuracy, there is still potential to refine the spatial–channel compression strategy. In future work, we aim to explore learnable token compression mechanisms that dynamically adapt to the content distribution in both spatial and channel dimensions. A promising approach is to integrate mask-guided autoregressive compression, inspired by MAE [he2022masked], to selectively retain informative tokens for global reasoning while discarding redundant ones. This adaptive compression could further reduce computational costs without sacrificing accuracy, enhancing the scalability and robustness of our framework for more complex dense prediction tasks.